# Is Watermarking LLM-Generated Code Robust?

### v1

@@ 1 Introduction
The rapid advancement of large language models (LLMs) like GPT and Codex (Chen et al., 2021; OpenAI, 2023) in understanding and generating code holds transformative potential for software engineering. However, it raises concerns about misuse such as code plagiarism and malware generation. Combating misuse requires accurate detection of LLM-generated code, which is challenging as LLMs are designed to produce realistic output that mimics human-generated code.
@@ A.4 Related Work
In practice, the text generated by a language model is likely modified by a user before being fed to a detector. These modifications may erase away the embedded hidden patterns and reduce the detectability of a watermark. As a result, a line of work has focused on designing robust watermarks for text that are detectable even if the original LLM output was changed. For instance, Zhao et al. (2023) simplify the soft watermarking scheme by consistently using a fixed red-green split and demonstrate that this new watermark is twice as robust to modifications as the baseline. Kirchenbauer et al discuss more robust detection schemes for when watermarked text is embedded in a larger human-written document. Additionally, Kuditipudi et al. (2023) propose a watermarking scheme that uses a key that is as large as the LLM-generated text and then aligns that key with the text to compute an alignment cost. They demonstrate that any text that is watermarked will be statistically likely to have low alignment costs and that this property holds even if the text was inserted or deleted from the original watermarked text. Recently, Christ et al. (2023) and Fairoze et al. (2023) proposed cryptographic watermarking schemes for text that achieve robustness properties.

### v2

@@ 1 Introduction
The rapid advancement of large language models (LLMs) like GPT and Codex in understanding and generating code holds transformative potential for software engineering (Chen et al., 2021; OpenAI, 2023; Ugare et al., 2024b). However, it raises concerns about misuse such as code plagiarism and malware generation. Combating misuse requires accurate detection of LLM-generated code, which is challenging as LLMs are designed to produce realistic output that mimics human-generated code.
@@ A.4 Related Work
In practice, the text generated by a language model is likely modified by a user before being fed to a detector. As a result, a line of work has focused on designing robust watermarks for text that are detectable even if the original LLM output was changed. For instance, Zhao et al. (2023) simplify the soft watermarking scheme by consistently using a fixed red-green split and demonstrate that this new watermark is twice as robust to modifications as the baseline. Kirchenbauer et al discuss more robust detection schemes for when watermarked text is embedded in a larger human-written document. Additionally, Kuditipudi et al. (2023) propose a watermarking scheme that uses a key that is as large as the LLM-generated text and then aligns that key with the text to compute an alignment cost. Recently, Christ et al. (2023) and Fairoze et al. (2023) proposed cryptographic watermarking schemes for text that achieve robustness properties.
@@
DNN Robustness A large body of research has focused on the robustness of LLMs and other DNNs against adversarial attacks (Zou et al., 2023; Ugare et al., 2024a; Zhang et al., 2023; Laurel et al., 2022; Ugare et al., 2023). This line of work is orthogonal to our investigation as we instead focus on the robustness of whether LLM-generated output (code) can be reliably detected.